OpenStudy (anonymous):

sqrt(80) - sqrt(125) + sqrt(45) Not really understanding how these problems work. The options to the question are: 5sqrt(2) -3sqrt(5) 3sqrt(5) 2sqrt(5) Any explanation/help would be appreciated :)

OpenStudy (anonymous):

@TheToneOfSurprise Here's how it works. :) First, you need to simplify all the terms. \[4\sqrt{5}-5\sqrt{5}+3\sqrt{5} \] Second, you need to add all the terms outside the square roots. (NOTE: Always remember to add or subtract only when the numbers inside the square roots are the same.) \[4\sqrt{5}-5\sqrt{5}+3\sqrt{5} = 2\sqrt{5} \] That's it! Hope it helps. :)

OpenStudy (anonymous):

I think I figured it out haha. I simplified each equation to 4sqrt(5) - 5sqrt(5) + 3sqrt(5) 4 -5 = -1 -1 + 3 = 2 Therefore, 2sqrt(5)?

OpenStudy (anonymous):

@TheToneOfSurprise That's great! :)

OpenStudy (anonymous):

@SweetsAndEverything Thanks for your help :) I awarded you best response or whatever - not sure what that does, but you deserve it either way haha.

OpenStudy (anonymous):

@TheToneOfSurprise I'm really not sure what that award is too but thank you. :)